OpenStudy (anonymous):

determine whether the function .on the given domain S is invertible. If it is , then find the inverse : f(x)= ln x S=[1,infinte) @ikram002p

OpenStudy (ikram002p):

its invertible when its continues among the given interval

OpenStudy (ikram002p):

so its invertible

OpenStudy (ikram002p):

y =ln x e^y= e^lnx e^y= x sup y=x & x=y the invers fun f^-1(x)=e^x among the interval [0,infinite)

OpenStudy (anonymous):

ok شكرررررررررررا

OpenStudy (anonymous):

سؤال تاني ممكن؟

OpenStudy (anonymous):

????????????

OpenStudy (ikram002p):

ok

OpenStudy (anonymous):

lim e^sinx =? x tend to 0

OpenStudy (ikram002p):

its continuse on x=0 so there is no problem wid the lmit oly apply x=0 lim e^sinx=e^sin 0 =1

OpenStudy (anonymous):

شكرا

OpenStudy (anonymous):

الاشتقاق y=arc sin 1/x find y'

OpenStudy (anonymous):

/??????????

OpenStudy (ikram002p):

is this online exam ?

OpenStudy (anonymous):

no i am sorry but i have a test near

OpenStudy (ikram002p):

oh ok.. cuz sry i dnt have time i have quiz to do good luck

OpenStudy (anonymous):

oky